The current in a wire is 5.0 amperes. Calculate the total amount of charge, in coulombs, that travels through the wire in 36 seconds.
1
Expert's answer
2020-06-11T10:33:28-0400

The current strength is a physical quantity equal to the ratio of the amount of charge ΔQ that has passed over the time Δt through the cross section of the conductor to the value of this period of time:


"I=\frac{\Delta Q}{\Delta t}=5 A"

the total amount of charge:


"\Delta Q=I\cdot\Delta t=5A \cdot 36sec= 180 C"

Answer: 180 C